### v1

@@ 4.1.2   Variational Autoencoder
$\log p_{\theta}(\mathbf{\hat{\mathbf{x}}}|\mathbf{\mathbf{x}})\propto\log e^{-|\mathbf{x}-\hat{\mathbf{x}}|^{2}}=-|\mathbf{x}-\hat{\mathbf{x}}|^{2},$ (4.3)
@@
$\begin{split}D_{KL}\left(q_{\phi}\left(\mathbf{z}|\mathbf{x}\right)||p(\mathbf{z})\right)&=D_{KL}\left(\mathcal{N}(\boldsymbol{\mu}_{1},\boldsymbol{\Sigma}_{1})||\mathcal{N}(\boldsymbol{\mu}_{2},\boldsymbol{\Sigma}_{2})\right)\\ &=\frac{1}{2}\left(\mathrm{Tr}(\boldsymbol{\Sigma}_{2}^{-1}\boldsymbol{\Sigma}_{1})+(\boldsymbol{\mu}_{2}-\boldsymbol{\mu}_{1})^{\top}\boldsymbol{\Sigma}_{2}^{-1}(\boldsymbol{\mu}_{2}-\boldsymbol{\mu}_{1})-D+\log\frac{\det\boldsymbol{\Sigma}_{2}}{\det\boldsymbol{\Sigma}_{1}}\right).\\ \end{split}$ (4.4)
@@
$\begin{split}D_{KL}\left(q_{\phi}\left(\mathbf{z}|\mathbf{x}\right)||p(\mathbf{z})\right)&=\frac{1}{2}\left(\mathrm{Tr}(\boldsymbol{\Sigma}_{1})+\boldsymbol{\mu}_{1}^{\top}\boldsymbol{\mu}_{1}-D-\log\det\boldsymbol{\Sigma}_{1}\right)\\ &=\frac{1}{2}\sum_{i=1}^{D}\left(\sigma_{1,i}^{2}+\mu_{1,i}^{2}-1-\log\sigma_{1,i}^{2}\right),\\ \end{split}$ (4.5)
where $\mu_{1,i}$ is $i$-th element in $\boldsymbol{\mu}_{1}$ and $\sigma_{1,i}$ is $i$-th element in $\boldsymbol{\sigma}_{1}$. The corresponding code implementation of the VAE loss is stated below. We use $\log\boldsymbol{\sigma}_{1}$ instead of $\boldsymbol{\sigma}_{1}$ for numerical stability, and the variable logvar corresponds to $2\log\boldsymbol{\sigma}_{1}$. The variable mu corresponds to $\boldsymbol{\mu}_{1}$. Figure 9 illustrates the structure of the neural network for this section. Figure 9 (a) shows the structure of the encoder which encodes the inputs to $D$-dimensional vectors for both Mean and Variance and we used $D=64$, and Figure 9 (b) shows the structure of the decoder which generates the sample from the latent vector. The code implementation is as follows:
@@
43 [...]
44 vae_model = VAE(input_dim=input_dim, hidden_dim=hidden_dim, z_dim=z_dim)
45
46 [...] # in training loop
47 x_reconstructed, mu, log_var = vae_model(x) # calls forward function in VAE class
48 recon_loss = F.mse_loss(x_reconstructed, x, reduction=’sum’) # Equation (4.3)
49 kl_loss = -0.5 * torch.sum(1 + logvar - mu.pow(2) - logvar.exp()) # Equation (4.4)
@@
4.1.3   Generative Adversarial Networks
Unlike VAEs, GANs do not require complex derivations of the loss function. The code implementation of the Equations (2.13) and (2.14) can be written intuitively. In practice, this involves using binary cross-entropy loss. The discriminator tries to identify if the given input is from the real data set or the generated data set. On the contrary, the generator tries to overcome the identification from the discriminator. Similar to the VAE, we constructed the 3-layer neural network for (a) generator and (b) discriminator as shown in Figure 10. The noise vector (latent vector) is $D$-dimensional vector where $D=64$. The code implementation is as follows:
@@ 4.1.3   Generative Adversarial Networks
30 [...]
31 generator = Generator(input_dim=z_dim, output_dim=input_dim, hidden_dim=hidden_dim)
32 discriminator = Discriminator(input_dim=input_dim, hidden_dim=hidden_dim)
33
34 criterion = nn.BCELoss()
35 real_label = torch.ones(batch_size)
36 fake_label = torch.zeros(batch_size)
37
38 [...] # in training loop
39 z = torch.randn(batch_size, z_dim) # random number generator from Standard Gaussian Distribution (N(0,1))
40 fake_data = generator(z) # Figure 10 (b)
@@
42 d_real = discriminator(real_data) # Figure 10 (a)
43 d_fake = discriminator(fake_data) # Figure 10 (a)
44 d_loss_real = criterion(d_real, real_label)
45 d_loss_fake = criterion(d_fake, fake_label)
46 [...] # update discriminator
47
48 z = torch.randn(batch_size, z_dim) # random number generator from Standard Gaussian Distribution (N(0,1))
49 fake_data = generator(z) # Figure 10 (b)
50 g_fake = discriminator(fake_data) # Figure 10 (a)
51 g_loss = criterion(discriminator(fake_data), real_label)
52 [...] # update generator
@@
4.1.4   Normalizing Flows (Flow-based Generative Models)
We use a flow-based generative model inspired by RealNVP (Dinh et al., 2016), which leverages affine coupling layers and alternating masking strategies to transform complex data distributions into a standard normal distribution. For simplicity, and since the data dimensionality is not too large, the model in this tutorial does not include permutation layers between coupling layers and batch normalization that were proposed in the original RealNVP paper. A more detailed implementation of RealNVP will be discussed in Section 4.2.4. Despite these simplifications, our model maintains the core idea of using invertible affine transformations to compute the log-likelihood and optimize the data distribution. The training objective is to maximize the likelihood of the observed data under this transformation, which can be expressed as minimizing the negative log-likelihood. The original loss function for the normalizing flow is the same as stated in Equation (2.21). For the affine coupling layers, we use the same equations as described from Equation (2.23) to Equation (2.26). We use much simpler networks for scaling and translation network as shown in Figure 11. The code implementation is as follows:
@@ 4.1.4   Normalizing Flows (Flow-based Generative Models)
8 y = x_masked + (1 - self.mask) * (x * torch.exp(s) + t) # Equation (2.23)
9 log_det_jacobian = torch.sum(s, dim=1) # Equation (2.26)
@@
33 log_prob = log_prob_z + log_det_jacobian # Equation (2.20)
@@
35 [...]
36 log_prob = flow_model(x)
37 loss = -torch.mean(log_prob) # Equation (2.21)
4.1.5   Diffusion Models
In the tutorial code, we used the Denoising Diffusion Probabilistic Model (DDPM) as the reference for generating data (Ho et al., 2020). Instead of using the traditional U-Net architecture for noise prediction, we implemented a simpler neural network structure, as shown in Figure 12. The U-Net architecture is commonly used in diffusion models due to its powerful ability to capture multi-scale features in image data. However, in this tutorial, we implemented a simpler neural network to reduce computational complexity and facilitate a focus on the fundamental concepts of diffusion models, such as stepwise noise estimation and progressive denoising. The loss function for training is calculated based on Equation (2.39) presented above, which measures the model’s ability to predict noise accurately at each step of the diffusion process. The code implementation is as follows:
Figure 12: Neural network structure of DDPM in HTS data
1 class DiffusionModel(nn.Module):
2 [...]
3
4 def forward_diffusion_sample(x0, t):
5 noise = torch.randn_like(x0)
6 sqrt_alpha_prod = sqrt_alphas_cumprod[t].view(-1, 1)
8 x_t = sqrt_alpha_prod * x0 + sqrt_one_minus_alpha_prod * noise
9 return x_t, noise
10 [...]
11 t = torch.randint(0, T, (batch_size,), device=device).long()
12 x_t, noise = forward_diffusion_sample(x0, t)
13 noise_pred = diffusion_model(x_t, t)
14 loss = nn.functional.mse_loss(noise_pred, noise) # Equation (2.39)
4.1.6   Score-based Generative Model
@@
$\begin{split}&\left|\left|{\nabla_{\mathbf{x}}\log q_{\sigma}\left(\tilde{\mathbf{x}}|\mathbf{x}\right)-s_{\theta}(\tilde{\mathbf{x}},\sigma)}\right|\right|^{2}_{2}=\left|\left|{\frac{\left(\tilde{\mathbf{x}}-\mathbf{x}\right)}{\sigma^{2}}+s_{\theta}(\tilde{\mathbf{x}},\sigma)}\right|\right|^{2}_{2},\\ \end{split}$ (4.6)
since,
$\begin{split}&q_{\sigma}\left(\tilde{\mathbf{x}}|\mathbf{x}\right)=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{\left(\tilde{\mathbf{x}}-\mathbf{x}\right)^{2}}{2\sigma^{2}}\right),\\ &\log q_{\sigma}\left(\tilde{\mathbf{x}}|\mathbf{x}\right)=-\frac{\left(\tilde{\mathbf{x}}-\mathbf{x}\right)^{2}}{2\sigma^{2}}-\log\left(\sqrt{2\pi\sigma^{2}}\right),\\ &\nabla_{\mathbf{x}}\log q_{\sigma}\left(\tilde{\mathbf{x}}|\mathbf{x}\right)=-\frac{\left(\tilde{\mathbf{x}}-\mathbf{x}\right)}{\sigma^{2}}.\end{split}$ (4.7)
where the term $\tilde{\mathbf{x}}-\mathbf{x}$ corresponds to the variable noise. We use a neural network with three linear layers and ReLU activations to obtain the estimated score as shown in Figure 13. In this tutorial, we train the NCSN using sequential sigma noise levels rather than randomly mixing them. This approach diverges from the typical method of random sampling of sigma values but was chosen to provide a clearer illustration of the overall training process. Given the relatively low dimension of the dataset, the sequential method is also feasible. The code implementation is as follows:
@@
32 loss = ((score + noise / (sigma ** 2)) ** 2).mean() # Equation (4.6)
@@
4.1.7   Results Comparison
Figure 14 presents the results of generating HTS data. In this section, we aim to share the empirical insights gained during the training of the models.
Similar to how VAEs tend to generate slightly blurry images in image generation, in this experiment with discrete values, the probability distribution also appeared relatively uniform rather than being overly skewed towards one side. However, it often struggled to accurately match the activity type and mode type, depending on the specified conditions. This may be attributed to the increased complexity resulting from the higher number of attributes.
In contrast, the GAN frequently encountered mode collapse, leading to cases where all probabilities were assigned to a particular activity or mode. However, this issue was less pronounced in cases involving fewer data attributes, such as problems related to origin location and destination location type.
The Normalizing Flow model demonstrated superior performance, with robustness across a wide range of hyperparameters. The loss reduction pattern followed the typical behavior expected in learning models, with an initial sharp decrease followed by a gradual decline in later stages. Throughout the training process and in the results, the model consistently performed well.
Figure 14: Results of each DGM in HTS data generation
The Diffusion model, which is commonly built using a U-Net architecture, was instead implemented with a simpler three-layer neural network in our approach. Despite initial concerns regarding its effectiveness, the model performed relatively well due to the simplicity of the dataset. However, occasional loss spikes occurred, which led the model to failure of training. Overall, the learning process remained stable, but future experiments could benefit from more sophisticated neural network architectures to further enhance performance.
The NCSN model exhibited high sensitivity to hyperparameters. For instance, when the max and min sigma values were low, the model displayed characteristics of high entropy or temperature, leading to unpredictable performance. Conversely, higher sigma values resulted in one-hot-like behavior. Additionally, the model’s performance varied across different categories: when it performed well on generating the location type, performance on generating types of action and mode became poor, and vice versa. Balancing hyperparameters effectively remains a key challenge for ensuring consistent performance across all attributes.
4.2 Generating Highway Traffic Speed Contour
The generation of highway traffic speed contours plays a crucial role in the analysis and visualization of traffic dynamics over both time and space. Traditionally, traffic flow has been represented using a time-space diagram of vehicles, which depicts the movement of vehicles along a road segment over time, providing insight into congestion patterns and traffic bottlenecks. However, converting this into a time-space speed contour offers several advantages. While traditional time-space diagrams focus on the position of vehicles at specific times, speed contours emphasize variations in traffic speed across both temporal and spatial dimensions, offering a clearer representation of traffic conditions such as slowdowns, jams, or free-flowing traffic across broader time periods. In this tutorial, we generated the traffic speed contour data based on the vehicle trajectory data.
Speed contours offer a continuous and intuitive visualization of traffic speeds across different locations and time intervals, making them particularly useful for traffic management and planning. They enable transportation professionals to quickly identify and quantify areas of concern, such as sections of highways prone to congestion or inconsistent traffic flow, thereby facilitating better decision-making. Moreover, these contours can serve as input for traffic forecasting models, aiding in the development of responsive traffic control systems, infrastructure improvements, and policy interventions aimed at reducing congestion and improving road safety.
4.2.1   Data and Preprocessing
Figure 15: Comparison of before and after the interpolation
In our tutorial section, we applied the DGMs to the Tennessee Department of Transportation’s I-24 Mobility Technology Interstate Observation Network (MOTION) data (Gloudemans et al., 2023) [^1] https://i24motion.org/data. I-24 MOTION data contains the vehicle trajectories in a four-mile section of I-24 in the Nashville-Davidson County Metropolitan area. It captures the trajectory in a high spatial and temporal resolution by using 294 ultra-high definition cameras and AI algorithms from Vanderbilt University. The AI algorithm transforms the raw videos into detailed 2-dimensional trajectories which is suitable to study vehicle behavior and traffic flow. The data collection occurred from Monday, November 21, 2022, to Friday, December 2, 2022, encompassing all eastbound and westbound lanes. The collection was done during peak morning hours, from 6 AM to 10 AM. This dataset is significant as it includes a variety of days as well as trajectories during an accident that occurred on November 21. However, it is important to note that when a single vehicle appears multiple times across different cameras, each instance is treated as a separate vehicle. Therefore, the whole trajectory will be sliced into multiple short sequences of trajectories. Consequently, this data needs a vehicle-matching algorithm to obtain the full trajectories of the single vehicle, but it can be effectively utilized for constructing time-space traffic contours without matching the vehicles.

### v3

@@ F.1 Variational Autoencoder
$\log p_{\theta}(\mathbf{\hat{\mathbf{x}}}|\mathbf{\mathbf{x}})\propto\log e^{-|\mathbf{x}-\hat{\mathbf{x}}|^{2}}=-|\mathbf{x}-\hat{\mathbf{x}}|^{2},$ (F.1)
@@
$\begin{split}D_{KL}\left(q_{\phi}\left(\mathbf{z}|\mathbf{x}\right)||p(\mathbf{z})\right)&=D_{KL}\left(\mathcal{N}(\boldsymbol{\mu}_{1},\boldsymbol{\Sigma}_{1})||\mathcal{N}(\boldsymbol{\mu}_{2},\boldsymbol{\Sigma}_{2})\right)\\ &=\frac{1}{2}\left(\mathrm{Tr}(\boldsymbol{\Sigma}_{2}^{-1}\boldsymbol{\Sigma}_{1})+(\boldsymbol{\mu}_{2}-\boldsymbol{\mu}_{1})^{\top}\boldsymbol{\Sigma}_{2}^{-1}(\boldsymbol{\mu}_{2}-\boldsymbol{\mu}_{1})-D+\log\frac{\det\boldsymbol{\Sigma}_{2}}{\det\boldsymbol{\Sigma}_{1}}\right).\\ \end{split}$ (F.2)
@@
$\begin{split}D_{KL}\left(q_{\phi}\left(\mathbf{z}|\mathbf{x}\right)||p(\mathbf{z})\right)&=\frac{1}{2}\left(\mathrm{Tr}(\boldsymbol{\Sigma}_{1})+\boldsymbol{\mu}_{1}^{\top}\boldsymbol{\mu}_{1}-D-\log\det\boldsymbol{\Sigma}_{1}\right)\\ &=\frac{1}{2}\sum_{i=1}^{D}\left(\sigma_{1,i}^{2}+\mu_{1,i}^{2}-1-\log\sigma_{1,i}^{2}\right),\\ \end{split}$ (F.3)
where $\mu_{1,i}$ is $i$-th element in $\boldsymbol{\mu}_{1}$ and $\sigma_{1,i}$ is $i$-th element in $\boldsymbol{\sigma}_{1}$. The corresponding code implementation of the VAE loss is stated below. We use $\log\boldsymbol{\sigma}_{1}$ instead of $\boldsymbol{\sigma}_{1}$ for numerical stability, and the variable logvar corresponds to $2\log\boldsymbol{\sigma}_{1}$. The variable mu corresponds to $\boldsymbol{\mu}_{1}$. Figure 13 illustrates the structure of the neural network for this section. Figure 13 (a) shows the structure of the encoder which encodes the inputs to $D$-dimensional vectors for both Mean and Variance and we used $D=64$, and Figure 13 (b) shows the structure of the decoder which generates the sample from the latent vector. The Pytorch code implementation is as follows:
@@
43 def compute_loss(x, x_reconstructed, mu, log_var):
44 """
45 MSE based reconstruction loss and KL divergence loss for VAE.
46 """
47 # Reconstruction loss
48 recon_loss = nn.functional.mse_loss(x_reconstructed, x, reduction=’sum’)
49 # KL Divergence
50 kl_loss = -0.5 * torch.sum(1 + log_var - mu.pow(2) - log_var.exp())
51 # total loss
@@
53 return loss
F.2 Generative Adversarial Networks
Unlike VAEs, GANs do not require complex derivations of the loss function. The code implementation of the Equations (2.8) and (2.9) can be written intuitively. In practice, this involves using binary cross-entropy loss. The discriminator tries to identify if the given input is from the real data set or the generated data set. On the contrary, the generator tries to overcome the identification from the discriminator. Similar to the VAE, we constructed the 3-layer neural network for (a) generator and (b) discriminator as shown in Figure 14. The noise vector (latent vector) is $D$-dimensional vector where $D=64$. The Pytorch code implementation is as follows:
@@ F.2 Generative Adversarial Networks
30 def compute_discriminator_loss(discriminator, real_data, fake_data, criterion, real_label, fake_label):
31 batch_size = real_data.size(0)
32 # loss for real data
33 real_targets = torch.full((batch_size, 1), real_label, device=real_data.device)
34 d_real = discriminator(real_data)
35 d_loss_real = criterion(d_real, real_targets)
36
37 # loss for generated data
38 fake_targets = torch.full((batch_size, 1), fake_label, device=real_data.device)
39 d_fake = discriminator(fake_data)
40 d_loss_fake = criterion(d_fake, fake_targets)
@@
42 d_loss = d_loss_real + d_loss_fake
43 return d_loss
44
45 def compute_generator_loss(discriminator, fake_data, criterion, real_label):
46 batch_size = fake_data.size(0)
47 real_targets = torch.full((batch_size, 1), real_label, device=fake_data.device)
48 d_fake = discriminator(fake_data)
49 g_loss = criterion(d_fake, real_targets)
50 return g_loss
@@
F.3 Normalizing Flows (Flow-based Generative Models)
We use a flow-based generative model inspired by RealNVP (Dinh et al., 2016), which leverages affine coupling layers and alternating masking strategies to transform complex data distributions into a standard normal distribution. For simplicity, and since the data dimensionality is not too large, the model in this tutorial does not include permutation layers between coupling layers and batch normalization that were proposed in the original RealNVP paper. A more detailed implementation of RealNVP will be discussed in Section F.3. Despite these simplifications, our model maintains the core idea of using invertible affine transformations to compute the log-likelihood and optimize the data distribution. The training objective is to maximize the likelihood of the observed data under this transformation, which can be expressed as minimizing the negative log-likelihood. The original loss function for the normalizing flow is the same as stated in Equation (2.14). For the affine coupling layers, we use the same equations as described from Equation (2.16) to Equation (2.19). We use much simpler networks for scaling and translation network as shown in Figure 15. The code Pytorch implementation is as follows:
@@ F.3 Normalizing Flows (Flow-based Generative Models)
8 y = x_masked + (1 - self.mask) * (x * torch.exp(s) + t)
9 log_det_jacobian = torch.sum(s, dim=1)
@@
33 log_prob = log_prob_z + log_det_jacobian
@@
35
36 def compute_nf_loss(model, x):
37 log_prob = model(x)
38 loss = -torch.mean(log_prob)
39 return loss
F.4 Score-based Generative Model
@@
$\begin{split}&\left|\left|{\nabla_{\mathbf{x}}\log q_{\sigma}\left(\tilde{\mathbf{x}}|\mathbf{x}\right)-s_{\theta}(\tilde{\mathbf{x}},\sigma)}\right|\right|^{2}_{2}=\left|\left|{\frac{\left(\tilde{\mathbf{x}}-\mathbf{x}\right)}{\sigma^{2}}+s_{\theta}(\tilde{\mathbf{x}},\sigma)}\right|\right|^{2}_{2},\\ \end{split}$ (F.4)
since
$\begin{split}&q_{\sigma}\left(\tilde{\mathbf{x}}|\mathbf{x}\right)=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{\left(\tilde{\mathbf{x}}-\mathbf{x}\right)^{2}}{2\sigma^{2}}\right),\\ &\log q_{\sigma}\left(\tilde{\mathbf{x}}|\mathbf{x}\right)=-\frac{\left(\tilde{\mathbf{x}}-\mathbf{x}\right)^{2}}{2\sigma^{2}}-\log\left(\sqrt{2\pi\sigma^{2}}\right),\\ &\nabla_{\mathbf{x}}\log q_{\sigma}\left(\tilde{\mathbf{x}}|\mathbf{x}\right)=-\frac{\left(\tilde{\mathbf{x}}-\mathbf{x}\right)}{\sigma^{2}}.\end{split}$ (F.5)
where the term $\tilde{\mathbf{x}}-\mathbf{x}$ corresponds to the variable noise. We use a neural network with three linear layers and ReLU activations to obtain the estimated score as shown in Figure 16. In this tutorial, we train the NCSN using sequential sigma noise levels rather than randomly mixing them. This approach diverges from the typical method of random sampling of sigma values but was chosen to provide a clearer illustration of the overall training process. Given the relatively low dimension of the dataset, the results show that the sequential method is also feasible. The Pytorch code implementation is as follows:
@@
32 loss = ((score + noise / (sigma ** 2)) ** 2).mean()
@@
F.5 Diffusion Models
In the tutorial code, we used the Denoising Diffusion Probabilistic Model (DDPM) as the reference for generating data (Ho et al., 2020). Instead of using the traditional U-Net architecture for noise prediction, we implemented a simpler neural network structure, as shown in Figure 17. The U-Net architecture is commonly used in diffusion models due to its powerful ability to capture multi-scale features in image data. However, in this tutorial, we implemented a simpler neural network to reduce computational complexity and facilitate a focus on the fundamental concepts of diffusion models, such as stepwise noise estimation and progressive denoising. The loss function for training is calculated based on Equation (2.32) presented above, which measures the model’s ability to predict noise accurately at each step of the diffusion process. The Pytorch code implementation is as follows:
Figure 17: Neural network structure of DDPM in HTS data
1 class DiffusionModel(nn.Module):
2 [...]
3
4 def forward_diffusion_sample(x0, t):
5 noise = torch.randn_like(x0)
6 sqrt_alpha_prod = sqrt_alphas_cumprod[t].view(-1, 1)
8 x_t = sqrt_alpha_prod * x0 + sqrt_one_minus_alpha_prod * noise
9 return x_t, noise
10 [...]
11
12 def compute_diffusion_loss(model, x0, t, sqrt_alphas_cumprod, sqrt_one_minus_alphas_cumprod):
13 """
14 Do a forward diffusion with the given x0 and timestep t,
15 compute the MSE between the noise predicted by the model and the actual noise.
16 """
17 x_t, noise = forward_diffusion_sample(x0, t, sqrt_alphas_cumprod, sqrt_one_minus_alphas_cumprod)
18 noise_pred = model(x_t, t)
19 loss = nn.functional.mse_loss(noise_pred, noise)
20 return loss
Appendix G Tutorials for Generating Highway Traffic Speed Contour
G.1 Variational Autoencoder
We utilize the result of Equation (F.3) to compute the loss of the VAE. The code structure and implementation of the loss function are consistent with the approach presented in Section F.1. Figure 18 illustrates the structure of the neural network for traffic speed contour generation. Figure 18 (a) illustrates the structure of the encoder. The encoder encodes the data to Mean and Variance with a $D$-dimensional vector. In this section, $D=64$ is applied. Figure 18 (a) is the structure of the decoder that generates the image from the latent vector.
G.2 Generative Adversarial Networks
The loss function and model architecture of the GAN for traffic speed contour estimation follows the implementation outlined in Section F.2. We employ binary cross-entropy loss for both the discriminator and the generator.
G.3 Normalizing Flows (Flow-based Generative Models)
As we discussed in Section 2.5, among various Flow-based Generative Models such as Dinh et al. (2014); Rezende and Mohamed (2015); Kingma and Dhariwal (2018); Durkan et al. (2019), we implemented RealNVP (Dinh et al., 2016) because of its ability to generate high-dimensional distribution due to its flexibility. As we can see in Equation (2.14), we need to calculate the $\log p_{0}(z_{0})$ in Equation (2.13) to get the likelihood of the data. In the implementation of RealNVP, we can simplify the log-likelihood $\log\left(p(\mathbf{z})\right)$ by assuming $z$ is sampled from the standard normal distribution. The other term in Equation (2.13), the summation of the determinant of the Jacobian, is simplified by applying a coupling layer as described through Equation (2.15) to Equation (2.19). Thus, the summation of the determinant of the Jacobian can be simplified as $\sum_{j}s(x_{1:d})_{j}$. In our code application, the regularization loss for batch normalization and scale parameters are applied following the contents in RealNVP (Dinh et al., 2016).
G.4 Score-based Generative Models
In the code implementation of NCSN in traffic speed contour, we had minor adjustments from the code from Section F.4. By definition, $\tilde{\mathbf{x}}$ is defined as $\tilde{\mathbf{x}}=\mathbf{x}+\sigma\cdot\epsilon$. Then $\nabla_{\mathbf{x}}\log q_{\sigma}\left(\tilde{\mathbf{x}}|\mathbf{x}\right)$ in Equation (F.5) is transformed as $\nicefrac{{\epsilon}}{{\sigma}}$. Practically, $\sigma^{2}$ is multiplied by the objective function as the weight of each stage. The detailed implementation can be found in the code.
G.5 Diffusion Models
The loss function of DDPM is simplified as the difference between the predicted noise and the real noise as stated in Equation (2.32). The criteria of the difference, i.e., the norm of difference, can be varied by the problem. Our code is also designed to select one of three types of loss, which is $L_{1}$, MSE loss, and Huber loss. For simplicity, we inserted the code block only including the MSE loss.
The neural network that is used for the DDPM is illustrated in Figure 22. Figure 22 (a), Figure 22 (b), and Figure 22 (c) show the structure of the Residual Block, Backbone Block, and U-Net of DDPM, respectively. The U-Net structure that is applied in the neural network in DDPM enables one to learn the information from high-level features or features from previous layers. The residual block contains the time embedding as an input, which identifies the stage of the current generation process. A similar mechanism can also be found in the Score-based model, as the score-based model creates an image using the scheduled $\sigma$. The neural network in DDPM also includes the attention mechanism that captures the relationship between features. The attention mechanism was applied inside the Backbone block, between the Residual block and the group normalization.
Appendix H Smoothing Process of Highway Traffic Speed Contour
The detailed process of interpolation, which is based on Edie’s definition of traffic flow dynamics, is outlined in Table 12.